OpenStudy (anonymous):

Jim Moore opened a new savings account. He deposited $12,000 at 12 percent compounded semiannually. At the start of the fourth year, Jim deposits an additional $50,000 that is also compounded semiannually at 12 percent. At the end of 6 years, the balance in Jim Moore's account is: (Use the tables in the handbook)

OpenStudy (anonymous):

=12000(1+.06)^6*2 +50,000(1+.06)^2*2 =12000*2.012+50000*1.2624 =24144+63120 =87264

OpenStudy (anonymous):

hope its right :)...please tell me the ans too :)